# Reduction of Cosymplectic groupoids by cosymplectic moment maps

### v2

@@ 4. Consequences of the main theorem
This section is devoted to applying Theorem 3.4 to specific situations concerning the multiplicativity of cosymplectic forms in a Lie groupoid. In particular, we are inspired by the methods presented in [6, 17] but in the context of cosymplectic structures. We also explore some of the main results in [15] when symmetries and reduction are considered.
@@ 4.1. Symmetries and reduction at the infinitesimal level
Under the existence of symmetries of a Lie algebroid of a cosymplectic groupoid, it is interesting to study how integration is related to reduction. Motivated by the results in [6], we study the reductions of the IM-forms associated with the cosymplectic forms.
We state the geometrical setup as follows. Let $M$ be a Poisson manifold. Consider a Lie algebroid $A\Rightarrow M$, endowed with an IM 2-form and an IM 1-form $(\mu,\nu)$ so that $A\cong T^{*}M\oplus\mathbb{R}$ as in (4)-(7). In addition, let $G$ be a Lie group acting on the manifold $M$. Suppose that this action lifts to an action on $A\Rightarrow M$ by Lie algebroid automorphisms, such that $\mu$ is $G$-equivariant and $\nu$ is $G$-invariant, i.e.
$\mu\circ\varphi_{g}^{A}=(d\varphi_{g^{-1}}^{M})^{*}\circ\mu,\qquad\nu\circ\varphi_{g}^{A}=\nu$
for all $g\in G$. The superindex notation refers to the space where $G$ acts.
Let $J_{can}:T^{*}M\to\mathfrak{g}^{*}$ be the canonical symplectic moment map of the cotangent lifting of the $G$-action on $M$ to $T^{*}M$. It is defined by $\langle J_{can}(\alpha),v\rangle=\alpha(v_{M})$ for any $v\in\mathfrak{g}$. Thus, we obtain a Lie algebroid morphism
[…]
where $\mathfrak{g}^{*}$ is thought as a Lie algebroid over a point with trivial Lie bracket. Note that, this map is $G$-equivariant because maps $\mu$ and $\mathcal{J}_{can}$ are $G$-equivariant. Moreover, $J$ is a Lie algebroid morphism (see [6]).
From now on we assume that $J^{-1}(0)\subset A$ is a subbundle. If $\mu$ has constant rank, the former condition is equivalent to $\mu(A)\cap J_{can}^{-1}(0)$ having constant rank. In [6], the authors prove that the subbundle $J^{-1}(0)\subset A$ is also a Lie subalgebroid. Furthermore, they show that the quotient by the action $G$ inherits a unique Lie algebroid structure. If the action $G\curvearrowright M$ is free and proper, then $M/G$ is a smooth manifold with surjective submersion quotient map $p:M\to M/G$. In addition, since $J^{-1}(0)$ is $G$-invariant, we can consider the vector bundle $A_{red}\rightarrow M/G$. It is defined as the unique vector bundle over $M/G$ such that the pullback bundle $p^{*}(A_{red})$ is naturally isomorphic to $J^{-1}(0)$, and the pullback of sections defines an identification of $\Gamma(A_{red})$ with the space of $G$-invariant sections of $\Gamma(J^{-1}(0)).$
Thus, the vector bundle $A_{red}$ inherits a unique Lie algebroid structure $(A_{red},[\cdot,\cdot]_{red},\rho_{red})$ so that:
  • The pullback of sections $p^{*}:\Gamma(A_{red})\to\Gamma(J^{-1}(0))$ preserves the Lie brackets.
  • For any $a\in\Gamma(A_{red})$ the vector field $\rho_{red}(a)$ and $\rho(p^{*}a)$ are $p$-related.
Let $\mu_{0}:J^{-1}(0)\to J_{can}^{-1}(0)$ and $\nu_{0}:J^{-1}(0)\to\mathbb{R}$ be the restrictions of the IM-forms by the inclusion map $J^{-1}(0)\to A$. Note that $\mu_{0}$ and $\nu_{0}$ are $G$-equivariant and $G$-invariant, respectively. Hence, it is possible to define the bundle maps $\mu_{red}:A_{red}\to T^{*}(M/G)$ and $\nu_{red}:A_{red}\to\mathbb{R}$ by satisfying the identities
$p^{*}(\mu_{red}(a))=\mu_{0}(p^{*}(a))\quad\text{ and }\quad p^{*}(\nu_{red}(a))=\nu_{0}(p^{*}(a)),$ (9)
where $a\in\Gamma(A_{red})$. The following diagrams illustrate the pullback maps that appear in these equations,
[…]
As presented below, a straightforward computation shows that the bundle maps $\mu_{red}$ and $\nu_{red}$ satisfy equations (4), (5), and (6), respectively.
@@
The forms $\mu_{red}$ and $\nu_{red}$ are IM-form on $A_{red}$.
@@
The skew-symmetry of $\mu_{red}$ follows directly from the definition. To check bracket compatibility, consider $a,b\in\Gamma(A_{red})$, and note that
$p^{*}\mu_{red}([a,b])$ $=\mu_{0}([p^{*}a,p^{*}b])=\mathcal{L}_{\rho(p^{*}a)}\mu_{0}(p^{*}b)-\imath_{\rho(p^{*}b)}d\mu_{0}(p^{*}a)$
$=\mathcal{L}_{\rho(p^{*}a)}p^{*}\mu_{red}(b)-\imath_{\rho(p^{*}b)}dp^{*}\mu_{red}(a)=p^{*}(\mathcal{L}_{\rho_{red}(a)}\mu_{red}(b)-\imath_{\rho_{red}(b)}d\mu_{red}(a)),$
and
@@
We now assume that $A$ is integrable by a source simply-connected groupoid. Let $(\mathcal{G},\omega,\eta)\rightrightarrows M$ be a cosymplectic groupoid, such that $Lie(\mathcal{G})=A$. The IM-forms $(\mu,\nu)$ defined by the forms $(\omega,\eta)$ satisfy (4)-(7) (see [15]). Let $G\curvearrowright A$ be an action as defined at the beginning of this section. From the Lie’s second theorem, the action $G\curvearrowright A$ integrates to an action $G\curvearrowright\mathcal{G}$ by Lie groupoid automorphisms. Notice that these two actions cover the $G$-action on $M$.
The Lie algebroid morphism $J:A\to\mathfrak{g}^{*}$ integrates to a $G$-equivariant Lie groupoid morphism
@@
Furthermore, it satisfies $\imath_{v_{\mathcal{G}}}\omega=-d\langle\mathcal{J},v\rangle$ for all $v\in\mathfrak{g}$ (see [6, Prop. 3.2]). Since $\mathcal{G}$ is a cosymplectic groupoid, the map $\mathcal{J}$ becomes a cosymplectic momentum map, as shown below.
@@
The action of $G$ on $\mathcal{G}$ is by cosymplectic diffeomorphism. Moreover, the map $-\mathcal{J}:\mathcal{G}\to\mathfrak{g}^{*}$ defined in (10) is a cosymplectic moment map, and $\xi\langle\mathcal{J},v\rangle=0$ for all $v\in\mathfrak{g}$.
@@
where $u\in A$ and $X\in TM$. Thus, from [5, Thm. 4.6] we conclude that $(\varphi_{g}^{\mathcal{G}})^{*}\omega=\omega$. Similarly, $G$-invariance of $\eta$ is proved by using that $\nu$ is $G$-invariant. To prove that $-\mathcal{J}$ is cosymplectic moment map we use that the $G$-action on $\mathcal{G}$ covers the action on $M$. In particular, we have $v_{\mathcal{G}}(x)=v_{M}(x)\in T_{x}M$ for $v\in\mathfrak{g}$ and $x\in M$. Because $\eta$ is multiplicative, then $\epsilon^{*}\eta=0$. Thus, $TM\subset\ker\eta$ and $(\imath_{v_{\mathcal{G}}}\eta)|_{M}=0$. Since $G$ preserves the form $\eta$, we deduce $0=\mathcal{L}_{v_{\mathcal{G}}}\eta=d\imath_{v_{\mathcal{G}}}\eta$. Therefore, $\imath_{v_{\mathcal{G}}}\eta=0.$ Furthermore, from $\imath_{v_{\mathcal{G}}}\omega=-d\langle\mathcal{J},v\rangle$, we conclude $-\xi\langle\mathcal{J},v\rangle=\imath_{\xi}\imath_{v_{\mathcal{G}}}\omega=0$. ∎
If $G$ acts freely on $M$, then by Lemma 3.3, 0 is a regular value of $\mathcal{J}$. Let $i:\mathcal{J}^{-1}(0)\to\mathcal{G}$ the inclusion map. Thus, the restricted source map satisfies $\ker(d(i^{*}s))_{g}=\ker(ds)_{g}\cap\ker(d\mathcal{J})_{g}$. Hence, we verify
$\ker(d(i^{*}s))|_{M}=A\cap\ker(d\mathcal{J})|_{M}=J^{-1}(0),$
making $\mathcal{J}^{-1}(0)$ an integration of $J^{-1}(0)$. If $G$ also acts properly on $M$, the quotient $\mathcal{G}_{red}=\mathcal{J}^{-1}(0)/G$ is a smooth manifold. We further recall that the action $G\curvearrowright\mathcal{G}$ is by groupoid automorphisms, so the quotient map $P:\mathcal{J}^{-1}(0)\to\mathcal{G}_{red}$ is a groupoid morphism. By differentiating $P$, we get $dP|_{J^{-1}(0)}:A_{red}\to Lie(\mathcal{G}_{red}).$ The kernel of $dP|_{J^{-1}(0)}$ is trivial, thus there is an identification between $A_{red}$ and $Lie(\mathcal{G}_{red})$ (see [6, Prop. 5.2]).
We summarize the results of this section in the following theorem, which is the main result of this section. In the next theorem, we suppose the source simply connected groupoid $\mathcal{G}$.
@@
Let $(\mathcal{G},\omega,\eta)\rightrightarrows M$ be a cosymplectic groupoid and $(A,\mu,\nu)$ its Lie algebroid. Let $G$ be a Lie group acting freely and properly on $M$. Suppose that this action lifts to an action on $A$ by Lie algebroid automorphisms, such that $\mu$ is $G$-equivariant and $\nu$ is $G$-invariant. Let $J$ be the map in (8), and $\mathcal{J}$ the map in (10). If $J^{-1}(0)\subset A$ is a subbundle, then the reduced Lie algebroid $(A_{red},\mu_{red},\nu_{red})$ is integrable by the cosymplectic groupoid $(\mathcal{G}_{red},\omega_{red},\eta_{red})\rightrightarrows M/G$.
@@
The algebroid $(A,\mu,\nu)$ satisfies the conditions of the geometrical setup established at the beginning of this section, where the Poisson structure on $M$ is given by $\pi_{M}=t_{*}\pi_{\mathcal{G}}$. By using the Lie algebroid morphism $J:A\to\mathfrak{g}^{*}$ defined in (8), we consider the algebroid $(A_{red},\mu_{red},\nu_{red})$ over $M/G$. Furthermore, the forms $\mu_{red}$ and $\nu_{red}$ are IM-forms by Lemma 4.1.
By Lie’s second theorem we integrate the $G$-action on $A$ to a $G$-action on $\mathcal{G}$, and we also integrate the map $J:A\to\mathfrak{g}^{*}$ to a groupoid morphism $\mathcal{J}:\mathcal{G}\to\mathfrak{g}^{*}$. Lemma 4.2 states that the map $\mathcal{J}$ is also a cosymplectic moment map. By using Theorem 3.4, we conclude that $(\mathcal{G}_{red},\omega_{red},\eta_{red})\rightrightarrows M/G$ is a cosymplectic Lie groupoid. As mentioned above, $A_{red}=Lie(\mathcal{G}_{red})$. Furthermore, let $u\in A_{red}$, then using the groupoid morphism $(P,p):(\mathcal{J}^{-1}(0)\to M)\to(\mathcal{G}_{red}\to M/G)$, we have
@@
In other words, for a cosymplectic groupoid the reduction of its infinitesimal counterpart is the same as the infinitesimal version of its cosymplectic reduction. Thus, we have that
${(\mathcal{G},\omega,\pi)\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$Liered${(\mathcal{G}_{red},\omega_{red},\pi_{red})\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$Lie${(A,\mu,\nu)\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$${(A_{red},\mu_{red},\nu_{red}),}$
@@
Let $(T^{*}M\oplus\mathbb{R},\mu=pr_{1},\nu=pr_{2})$ be a central extension Lie algebroid with brackets as in (7). If there is a $G$ action on $M$ so that lifts to action by Lie algebroid morphism on $A$ so that $\mu$ is equivariant and $\nu$ is invariant, then
$(T^{*}(M/G)\oplus\mathbb{R},\mu_{red}=pr_{1},\nu_{red}=pr_{2})$
@@ 4.2. Reduction of the symplectic leaf
Given a cosymplectic groupoid $(\mathcal{G},\omega,\eta)\rightrightarrows M$, let $\Sigma$ be the symplectic leaf of $\pi_{\mathcal{G}}$ containing $M$. Let $L:\Sigma\to\mathcal{G}$ denote the inclusion. In [15] it is proved that $\Sigma\subset\mathcal{G}$ is a Lie subgroupoid, and $(\Sigma,L^{*}\omega)\rightrightarrows M$ is a symplectic groupoid integrating $\pi_{M}$. Assuming that we impose symmetries on the cosymplectic groupoid $(\mathcal{G},\omega,\eta)$ so that we can construct its cosymplectic reduced groupoid, we inquire about the symmetries and the reduction of $(\Sigma,L^{*}\omega)\rightrightarrows M$. We want to restrict the action $G$ to the leaf $\Sigma$, and we want to define a symplectic moment map.
To show that the action of $G$ on $\mathcal{G}$ is restricted to an action on $\Sigma$, we first state a basic fact about morphisms between Poisson varieties.
@@
Let $M$ and $N$ be Poisson manifold, and $\varphi:M\to N$ a Poisson diffeomorphism. Then, $\varphi$ preserves the symplectic foliation.
@@
Because $\varphi$ is a Poisson morphism, the Hamiltonian vectors are $\varphi$-related. Since the symplectic foliations in $M$ and $N$ are generated by their respective Hamiltonian vector fields, then the leaves in $M$ correspond to the leaves in $N$. In fact, let $f\in C^{\infty}(M)$ and $p,p^{\prime}\in M$ be on a same leaf. Let $\gamma$ be an integral curve of $X_{f}\in\mathfrak{X}(M)$ joining $p$ and $p^{\prime}$, therefore the curve $\tilde{\gamma}=\varphi\circ\gamma$ is the solution of $\varphi_{*}X_{f}\in\mathfrak{X}(N)$ joining $\varphi(p)$ and $\varphi(p^{\prime})$. ∎
It is important to highlight that the leaves of the symplectic foliation may be non-invariant by the action of a group by Poisson automorphisms (cf. [24, Ex. 4.2.5]). Moreover, in the case of standard momentum maps on Poisson manifolds, a necessary condition for the existence of a momentum map is that the action by Poisson morphisms being leaf preserving, which does not happen in general [24, §4.5]. However, in the case of a cosymplectic groupoid, the action does preserve the symplectic leaf containing the units. Thus, we can restrict the action to this leaf.
@@
The action $G\curvearrowright\mathcal{G}$ is by Poisson diffeomorphisms. By Lemma 4.5, any left multiplication either transforms the leaf $\Sigma$ into another leaf of the symplectic foliation, or leaves it invariant. Furthermore, for any $g\in G$, the units $M$ are contained in $\varphi_{g}(\Sigma)$. Therefore, by the definition of $\Sigma$, we have $\varphi_{g}(\Sigma)=\Sigma$. In addition we verify that
@@
[…]
@@
Let $G$ be a Lie group acting on a cosymplectic groupoid $(\mathcal{G},\omega,\eta)\rightrightarrows M$, and $\mathcal{J}:\mathcal{G}\to\mathfrak{g}^{*}$ a cosymplectic moment map satisfying the conditions of Theorem 3.4. Let $\Sigma$ the symplectic leaf containing $M$. Then the groupoid morphism $\mathcal{J}_{0}:(\Sigma,L^{*}\omega)\to\mathfrak{g}^{*}$ defined by $L^{*}\mathcal{J}$ is a symplectic moment map with respect to the $G$-action in Lemma 4.6.
@@
Since the action $G$ on $\Sigma$ is the restriction of the action $G\curvearrowright\mathcal{G}$, for any $g\in G$ the left multiplication satisfies $L\circ\varphi_{g}=\varphi_{g}\circ L$. Thus, the $G$-equivariance of $\mathcal{J}_{0}$ follows from the $G$-equivariance of $\mathcal{J}$, that is $Ad_{g}^{*}\mathcal{J}_{0}=\mathcal{J}\circ\varphi_{g}\circ L=\mathcal{J}_{0}\varphi_{g}.$
Moreover, for $v\in\mathfrak{g}$, its fundamental vector field $v_{\mathcal{G}}|_{\Sigma}$ is actually a vector field on $\Sigma$. Hence, $L^{*}\imath_{v_{\mathcal{G}}}\omega=\imath_{v_{\mathcal{G}}}L^{*}\omega$. Since $\mathcal{J}$ is a cosymplectic moment map and $\xi(\langle J,v\rangle)=0$, we conclude

### v3

@@ 4. Consequences of the main theorem
This section is devoted to applying Theorem 3.4 to specific situations concerning the multiplicativity of cosymplectic forms in a Lie groupoid. In particular, we are inspired by the methods presented in [BC2, FOR] but in the context of cosymplectic structures. We also explore some of the main results in [FI] when symmetries and reduction are considered.
@@ 4.1. Symmetries and reduction at the infinitesimal level
Under the existence of symmetries of a Lie algebroid of a cosymplectic groupoid, it is interesting to study how integration is related to reduction. Motivated by the results in [BC2], we study the reductions of the IM-forms associated with the cosymplectic forms.
We state the geometric setup for IMcs-algebroids as follows.
  • i)
    Let $(A,\mu,\nu,\zeta)$ be an IMcs algebroid over a Poisson manifold $(M,\pi_{M})$ (cf. Definition 2.16);
  • ii)
    a Lie group $G$ with an action $\varphi^{A}:G\times A\to A$ by Lie algebroid morphisms, covering an action on $M$ (denoted by $\varphi^{M}$);
  • iii)
    $\mu$ and $\zeta$ are $G$-equivariant and $\nu$ is $G$-invariant, i.e.
    $\mu\circ\varphi_{g}^{A}=(d\varphi_{g^{-1}}^{M})^{*}\circ\mu,\qquad\zeta\circ\varphi_{g}^{M}=\varphi_{g}^{A}\circ\zeta\qquad\nu\circ\varphi_{g}^{A}=\nu,\qquad\text{ for all }g\in G.$
Recall the map $J:A\to\mathfrak{g}^{*}$ from Proposition 2.7, and assume that $J^{-1}(0)$ is a vector subbundle of $A$ (for instance, if $\mu(A)$ and $\mu(A)\cap J_{can}^{-1}(0)$ have constant rank). Analogously to the construction of $\mu_{red}$ in Proposition 2.8, we obtain $\nu_{0}:J^{-1}(0)\to\mathbb{R}$ as the restriction via the inclusion map $J^{-1}(0)\to A$. Moreover, by the geometric setup, $\nu_{0}$ is $G$-invariant. Hence, it is possible to define the bundle map $\nu_{red}:A_{red}\to\mathbb{R}$ such that
$p^{*}(\nu_{red}(a))=\nu_{0}(p^{*}(a)),\qquad\text{ for all }a\in\Gamma(A_{red}).$ (10)
The following diagram illustrates the pullback map that appears in this equation,
${J^{-1}(0)\simeq p^{*}A_{red}\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$$\scriptstyle{\nu_{0}}$${A_{red}\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$$\scriptstyle{\nu_{red}}\textstyle{\mathbb{R}\ignorespaces\ignorespaces\ignorespaces\ignorespaces}\textstyle{\mathbb{R}}$${M\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$$\scriptstyle{p}$${M/G.}$
@@
The forms $\mu_{red}$ and $\nu_{red}$ are IM-forms on $A_{red}$, and there exists a section $\zeta_{red}\in\Gamma(A_{red})$ such that $\mathrm{Ker}(\mu_{red})=\langle\zeta_{red}\rangle.$
@@
The claim about $\mu_{red}$ has already been proved in Proposition 2.8, and the argument is analogous for $\nu$. For the reader’s convenience, we present the proof of (7) for $\nu_{red}$. Let $a,b\in\Gamma(A_{red})$. Then
@@
Finally, note that $\zeta\in\Gamma(A)$ belongs to $\Gamma(J^{-1}(0))$, since $\mu(\zeta)=0$ and $J=J_{can}\circ\mu$. Together with the $G$-equivariance of $\zeta$, this yields $\zeta_{red}\in\Gamma(A_{red})$ via the quotient map on sections $p^{*}:\Gamma(A_{red})\to\Gamma(J^{-1}(0))$. It remains to show that $\mathrm{Ker}(\mu_{red})=\langle\zeta_{red}\rangle$. Indeed, let $a_{r}\in\mathrm{Ker}(\mu_{red})$. Then
$0=p^{*}(\mu_{red}(a_{r}))=\mu_{0}(p^{*}a_{r})=\mu(a),$
where the second identity follows from the pullback bundle map $p^{*}$ and the last equality holds since $a\in\Gamma(J^{-1}(0))$. As $\mathrm{Ker}(\mu)=\langle\zeta\rangle$ and $\zeta\in\Gamma(J^{-1}(0))$, this implies $a=\zeta$, and therefore its quotient section is $a_{r}=\zeta_{red}$. ∎
We now assume that the IMcs-algebroid $(A,\mu,\nu,\zeta)$ (over a Poisson manifold $(M,\pi_{M})$) is integrated by a source-simply-connected cosymplectic Lie groupoid $(\mathcal{G},\omega,\eta,\xi)\rightrightarrows M$. Also consider the $G$-equivariant Lie groupoid morphism
@@
that satisfies $\imath_{v_{\mathcal{G}}}\omega=-d\langle\mathcal{J},v\rangle$ for all $v\in\mathfrak{g}$ (see Proposition 2.9). Since $\mathcal{G}$ is a cosymplectic groupoid, the map $\mathcal{J}$ becomes a cosymplectic momentum map, as shown below.
@@
The action of $G$ on $\mathcal{G}$ is by cosymplectic diffeomorphisms. Moreover, the map $\mathcal{J}:\mathcal{G}\to\mathfrak{g}^{*}$ defined in (11) is a cosymplectic moment map, and $\xi\langle\mathcal{J},v\rangle=0$ for all $v\in\mathfrak{g}$.
@@
where $u\in A$ and $X\in TM$. Thus, from [BC, Thm. 4.6] we conclude that $(\varphi_{g}^{\mathcal{G}})^{*}\omega=\omega$. Similarly, $G$-invariance of $\eta$ is proved by using that $\nu$ is $G$-invariant. To prove that $\mathcal{J}$ is a cosymplectic moment map we use that the $G$-action on $\mathcal{G}$ covers the action on $M$. In particular, we have $v_{\mathcal{G}}(x)=v_{M}(x)\in T_{x}M$ for $v\in\mathfrak{g}$ and $x\in M$. Because $\eta$ is multiplicative, then $\epsilon^{*}\eta=0$. Thus, $TM\subset\ker\eta$ and $(\imath_{v_{\mathcal{G}}}\eta)|_{M}=0$. Since $G$ preserves the form $\eta$, we deduce $0=\mathcal{L}_{v_{\mathcal{G}}}\eta=d\imath_{v_{\mathcal{G}}}\eta$. Therefore, $\imath_{v_{\mathcal{G}}}\eta=0.$ Furthermore, from $\imath_{v_{\mathcal{G}}}\omega=-d\langle\mathcal{J},v\rangle$, we conclude $\xi\langle\mathcal{J},v\rangle=\imath_{\xi}\imath_{v_{\mathcal{G}}}\omega=0$. ∎
As a first application of Theorem 3.4, and applying Proposition 2.9 together with the preceding lemmas, we obtain the main result of this section.
@@
Let $(A,\mu,\nu,\zeta)$ be an IMcs algebroid over a Poisson manifold $(M,\pi_{M})$, and let $G$ be a Lie group acting freely and properly on $M$. Assume that the lifted action of $G$ on $A$ preserves the IMcs-algebroid structure, as described in the geometric setup i)i)$iii)$. Consider the map $J:A\to\mathfrak{g}^{*}$ from Proposition 2.7. If $J^{-1}(0)$ is a vector subbundle of $A$, then:
  1. (1)
    $(A_{red},\mu_{red},\nu_{red},\zeta_{red})$ is an IMcs-algebroid over $(M/G,\pi_{red})$.
  2. (2)
    If $(\mathcal{G},\omega,\eta)\rightrightarrows M$ is the source simply-connected cosymplectic Lie groupoid integrating $A$, then $(\mathcal{G}_{red},\omega_{red},\eta_{red})\rightrightarrows M/G$ is a cosymplectic Lie groupoid integrating $(A_{red},\mu_{red},\nu_{red},\zeta_{red})$.
@@
For the first item, by Lemma 4.1, it suffices to check that $\mu_{red}:A_{red}\to T^{*}(M/G)$ is surjective. Take $\alpha_{r}\in T^{*}(M/G)$ and $a\in A$ such that $\mu(a)=p^{*}\alpha_{r}$. Then, for any $u\in\mathfrak{g}$,
$J(a)(u)=J_{can}(p^{*}\alpha_{r})(u)=i_{u_{M}}p^{*}\alpha_{r}=0.$
Hence, for every $\alpha_{r}\in T^{*}(M/G)$ there exists $a\in J^{-1}(0)$ with $\mu_{0}(a)=\alpha_{r}$. Passing to the quotient of sections, we obtain $a_{red}$ with $\mu_{red}(a_{red})=\alpha_{r}$. Thus, $\mu_{red}$ is a surjective bundle map, and yields the central extension
$0\to\langle\zeta_{red}\rangle\to A_{red}\overset{\mu_{red}}{\longrightarrow}T^{*}(M/G)\to 0,$
with $\underline{\nu}_{red}(a_{red})=\nu_{red}(a_{red})\zeta_{red}$ as a splitting.
For the second item, by Lie’s second theorem we integrate the $G$-action on $A$ to a $G$-action on $\mathcal{G}$, and we also integrate the map $J:A\to\mathfrak{g}^{*}$ to a groupoid morphism $\mathcal{J}:\mathcal{G}\to\mathfrak{g}^{*}$, analogously to Proposition 2.9. Lemma 4.2 states that the map $\mathcal{J}$ is also a cosymplectic moment map. By using Theorem 3.4, we conclude that $(\mathcal{G}_{red},\omega_{red},\eta_{red})\rightrightarrows M/G$ is a cosymplectic Lie groupoid. As mentioned in Proposition 2.9, $A_{red}=Lie(\mathcal{G}_{red})$. Furthermore, let $u\in A_{red}$, then using the groupoid morphism $(P,p):(\mathcal{J}^{-1}(0)\to M)\to(\mathcal{G}_{red}\to M/G)$, we have
@@
In other words, for a cosymplectic groupoid, the reduction of its infinitesimal counterpart is the same as the infinitesimal version of its cosymplectic reduction. Thus, we have that
[…]
@@
Let $(T^{*}M\oplus\mathbb{R},\mu=pr_{1},\nu=pr_{2},\zeta=\partial r)$ be a trivial central extension Lie algebroid with brackets as in (9). If there is a $G$ action on $M$ which lifts to an action by Lie algebroid morphism on $A$ so that $\mu$ is equivariant and $\nu$ is invariant, then
$(T^{*}(M/G)\oplus\mathbb{R},\mu_{red}=pr_{1},\nu_{red}=pr_{2},\partial r)$
@@ 4.2. Reduction of the symplectic leaf
Given a cosymplectic groupoid $(\mathcal{G},\omega,\eta)\rightrightarrows M$, let $\Sigma$ be the symplectic leaf of $\pi_{\mathcal{G}}$ containing $M$. Let $L:\Sigma\to\mathcal{G}$ denote the inclusion. In [FI] it is proved that $\Sigma\subset\mathcal{G}$ is a Lie subgroupoid, and $(\Sigma,L^{*}\omega)\rightrightarrows M$ is a symplectic groupoid integrating $\pi_{M}$. Assuming that we impose symmetries on the cosymplectic groupoid $(\mathcal{G},\omega,\eta)$ so that we can construct its cosymplectic reduced groupoid, we inquire about the symmetries and the reduction of $(\Sigma,L^{*}\omega)\rightrightarrows M$. We want to restrict the action $G$ to the leaf $\Sigma$, and we want to define a symplectic moment map.
To show that the action of $G$ on $\mathcal{G}$ is restricted to an action on $\Sigma$, we first state a basic fact about morphisms between Poisson manifolds.
@@
Let $M$ and $N$ be Poisson manifold, and let $\varphi:M\to N$ be a Poisson diffeomorphism. Then, $\varphi$ preserves the symplectic foliation.
@@
This fact follows from the $\varphi$-related Hamiltonian vector fields on the respective Poisson manifolds. ∎
It is important to highlight that the leaves of the symplectic foliation may be non-invariant by the action of a group by Poisson automorphisms (cf. [ORbook, Ex. 4.2.5]). Moreover, in the case of standard momentum maps on Poisson manifolds, a necessary condition for the existence of a momentum map is that the action by Poisson morphisms being leaf preserving, which does not happen in general [ORbook, §4.5]. However, in the case of a cosymplectic groupoid, the action does preserve the symplectic leaf containing the units. Thus, we can restrict the action to this leaf.
@@
The action $G\curvearrowright\mathcal{G}$ is by Poisson diffeomorphisms. By Lemma 4.5, any left multiplication either transforms the leaf $\Sigma$ into another leaf of the symplectic foliation, or leaves it invariant. Furthermore, for any $g\in G$, the units $M$ are contained in $\varphi_{g}(\Sigma)$. Therefore, by the definition of $\Sigma$, we have $\varphi^{\mathcal{G}}_{g}(\Sigma)=\Sigma$. Let $\varphi^{\Sigma}$ denote the action of $G$ on $\Sigma$. In addition, we verify that
@@
With this action of $G$ on $\Sigma$ by symplectic groupoid automorphisms, we can show that the groupoid morphism $\mathcal{J}_{0}:\Sigma\xrightarrow{L}\mathcal{G}\xrightarrow{\mathcal{J}}\mathfrak{g}^{*}$ is a symplectic moment map. For the convenience of the reader, we recall the assumptions of Theorem 3.4: Let $G$ be a Lie group acting on a cosymplectic groupoid $(\mathcal{G},\omega,\eta)\rightrightarrows M$ by cosymplectic groupoid automorphisms. Let $\mathcal{J}:\mathcal{G}\to\mathfrak{g}^{*}$ be a cosymplectic moment map such that it is a Lie groupoid morphism, and $\xi(\langle\mathcal{J},v\rangle)=0$ for all $v\in\mathfrak{g}$. Moreover, assume that $G$ acts freely and properly on $M$.
@@
Let $G$ be a Lie group acting on a cosymplectic groupoid $(\mathcal{G},\omega,\eta)\rightrightarrows M$, and let $\mathcal{J}:\mathcal{G}\to\mathfrak{g}^{*}$ be a cosymplectic moment map satisfying the conditions of Theorem 3.4. Let $\Sigma$ be the symplectic leaf containing $M$. Then the groupoid morphism $\mathcal{J}_{0}:(\Sigma,L^{*}\omega)\to\mathfrak{g}^{*}$ defined by $L^{*}\mathcal{J}$ is a symplectic moment map with respect to the $G$-action in Lemma 4.6.
@@
Since the action of $G$ on $\Sigma$ is the restriction of the action $\varphi^{\mathcal{G}}:G\curvearrowright\mathcal{G}$, for any $g\in G$ $L\circ\varphi^{\Sigma}_{g}=\varphi^{\mathcal{G}}_{g}\circ L$. Thus, the $G$-equivariance of $\mathcal{J}_{0}$ follows from the $G$-equivariance of $\mathcal{J}$, that is $Ad_{g}^{*}\mathcal{J}_{0}=\mathcal{J}\circ\varphi^{\mathcal{G}}_{g}\circ L=\mathcal{J}_{0}\varphi^{\Sigma}_{g}.$
Moreover, for $v\in\mathfrak{g}$, the restriction $v_{\mathcal{G}}|_{\Sigma}$ of its fundamental vector field is actually a vector field on $\Sigma$. Hence, $L^{*}\imath_{v_{\mathcal{G}}}\omega=\imath_{v_{\mathcal{G}}}L^{*}\omega$. Since $\mathcal{J}$ is a cosymplectic moment map and $\xi(\langle J,v\rangle)=0$, we conclude